EliDaProfit:

-2x5-7 what's the total

XxXNessalulbaddieXxX:

do u know what u do first?

EliDaProfit:

yes

XxXNessalulbaddieXxX:

@elidaprofit wrote:
yes
mk what do u do first

EliDaProfit:

you flip the sighns

XxXNessalulbaddieXxX:

@elidaprofit wrote:
you flip the sighns
well go ahead and do that

EliDaProfit:

i did but i still don't understand

XxXNessalulbaddieXxX:

@elidaprofit wrote:
i did but i still don't understand
hmmm what part do u not understand :thonk:

Florisalreadytaken:

can you please provide a ss of the question? because im unsure if its \( -2x \times5-7 \) or \(-2 \times5-7 \)

Yugiv8:

maybe the 2x5-7= 3

jhonyy9:

i think -2×5 -7 = ? make sense

Florisalreadytaken:

i ws just making sure, as if it is the 1st one, we will have a little more stepsthan the 2nd one as we had to equalise it all by 0 and blla blla-- butlets go with the 2nd one haha so we have got \(-2 \times5-7\) we first have to do the multiplication, so it will look like this \[ (-2 \times5)-7 \] what is negative two times 5?

Florisalreadytaken:

NOTE: \( (-)\times(+)= (-) \)

EliDaProfit:

I don't know

EliDaProfit:

-2x5=-10

Florisalreadytaken:

its simple! we said that \( (-)\times(+)= (-) ​\) now lets just multiply the numbers as we would usually do: \( 2\times5=10 \) adding the negative sign, we get \( -10 \) so we're left with: \( -10-7=x ​\) that is the same as writing \( (-10)+(-7)=x \) what is x?

EliDaProfit:

-17

EliDaProfit:

-17

EliDaProfit:

-17

Yugiv8:

-10?

Florisalreadytaken:

its negative 17, you are correct!